OpenStudy (anonymous):

Factor Completely: x^2 - (y - 6)^2

OpenStudy (kinggeorge):

Notice that this is a difference of 2 squares. In general, \[x^2-a^2=(x-a)(x+a)\]Here, you have \(a=y-6\). Can you tell me what the factorization is now?

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

where are you getting the a from?

OpenStudy (kinggeorge):

The \(a\) is just some random variable I used. Since \(x^2-a^2=(x-a)(x+a)\), and \(a=y-6\) using a simple substitution you get that \[x^2-(y-6)^2=(x-(y-6))(x+(y-6))\]

OpenStudy (anonymous):

oooooohhhhh okay I get it!

OpenStudy (kinggeorge):

:)

OpenStudy (anonymous):

thank you!

OpenStudy (kinggeorge):

You're welcome.